### v1

# Hints of Prompt: Enhancing Visual Representation for Multimodal LLMs in Autonomous Driving

Abstract
In light of the dynamic nature of autonomous driving environments and stringent safety requirements, general MLLMs combined with CLIP alone often struggle to represent driving-specific scenarios accurately, particularly in complex interactions and long-tail cases. To address this, we propose the Hints of Prompt (HoP) framework, which introduces three key enhancements: Affinity hint to emphasize instance-level structure by strengthening token-wise connections, Semantic hint to incorporate high-level information relevant to driving-specific cases, such as complex interactions among vehicles and traffic signs, and Question hint to align visual features with the query context, focusing on question-relevant regions. These hints are fused through a Hint Fusion module, enriching visual representations and enhancing multimodal reasoning for autonomous driving VQA tasks. Extensive experiments confirm the effectiveness of the HoP framework, showing it significantly outperforms previous state-of-the-art methods across all key metrics.
00footnotetext: Work conducted as an intern at DeepRoute.AI. †Corresponding author.
Figure 1: Overview of the Hints of Prompt (HoP) Framework. HoP enhances Visual Question Answering (VQA) in autonomous driving by incorporating three hierarchical hints: Affinity, Semantic, and Question. The Affinity hint provides foundational instance-level structures through token-wise connections, aiding in instance boundary and interaction recognition. Building on this, the Semantic hint introduces specific instances along with their category information, adding essential driving-related contexts, such as vehicles and traffic signs. Finally, the Question hint guides the LLM’s attention toward image regions pertinent to the question. These hints are fused with visual tokens through a simple Hint Fusion module, aligned via an adapter, and then processed by the LLM to generate answers.
footnotetext: * Work conducted as intern at DeepRoute.AI. $\dagger$ Corresponding author.
1 Introduction
In recent years, there have been growing trends towards formalizing autonomous driving within end-to-end learning-based frameworks [18]. Leveraging extensive human driving data, these approaches demonstrate human-like behavior and generalization capabilities, enabling them to handle intricate scenarios more effectively. However, the end-to-end pipeline often functions as a black box, making it challenging to interpret or explain the neural network-driven planning processes [7]. This opacity raises significant ethical and legal concerns regarding accountability and trustworthiness in safety-critical applications.
Figure 2: Visualization of token affinity from CLIP [37] and DINOv2 [35]. Similar colors indicate higher affinity scores, with color values derived from a PCA-reduced token vector space. DINOv2 token-wise similarity denotes tokens embedded only from the similarity matrix.
Interpretability has emerged as a crucial research focus in autonomous driving to address these concerns. Visual Question Answering (VQA) tasks have gained considerable attention in this context [39], as they allow systems to articulate their observations and actions through natural language interactions. By bridging the gap between complex decision-making processes and user comprehension, VQA enhances system transparency and fosters user trust.
The application of Multimodal Large Language Models (MLLMs) [4, 46, 42, 8, 26, 14] has significantly advanced VQA tasks by integrating visual encoders with Large Language Models (LLMs) [12, 43, 10] for deeper multimodal understanding. Typically, MLLMs comprise a visual encoder, an adapter, and an LLM. For VQA, the visual encoder extracts image features, aligned by the adapter with text tokens. The LLM tokenizes the question and integrates the aligned visual tokens to generate an answer.
For VQA in autonomous driving, the dynamic nature of driving environments and stringent safety requirements pose challenges for general MLLMs, particularly in representing hierarchical visual features crucial for complex interactions and long-tail cases. Although CLIP-based models [37, 50] have dominated visual encoders, as shown in Fig. 2, CLIP visual tokens overlook token-wise affinity, which is essential for preserving instance-level structure and establishing basic objectness understanding in driving scenarios. In contrast, DINOv2 [35] better maintains this structure. Additionally, general visual encoders underrepresent domain-specific semantics, making them prone to missing small yet crucial elements for driving, such as distant vehicles, pedestrians, and traffic signs—features vital for reasoning in complex interactions, like pedestrians unexpectedly crossing the street. Lastly, current MLLMs struggle to adjust visual representations based on specific questions, as visual and text tokens are processed separately. This separation limits the model’s ability to focus on query-relevant image regions, reducing the effectiveness of context-aware responses in autonomous driving.
Based on these observations, we introduce stratified tokens as hints to enhance visual representation. Specifically, we propose three key enhancements: Affinity hint to capture instance-level structure and strengthen token-wise connections, Semantic hint to incorporate high-level semantic information for domain-specific elements, and Question hint to correlate visual features according to the query context, ensuring focus on question-relevant regions. We conduct comprehensive studies on designing effective and efficient hints fusion strategies for combining question-specific and hierarchical semantic cues tailored to autonomous driving, thereby enhancing multimodal reasoning in these scenarios. Moreover, we propose an efficient version of HoP that maintains competitive performance, achieved by extracting Affinity and Semantic hints directly from two lightweight heads distilled and fine-tuned from the CLIP backbone. Fig. 1 illustrates our proposed method, referred to as Hints of Prompt (HoP) framework throughout the paper.
Extensive experiments demonstrate significant improvements for VQA. On the LingoQA [34] benchmark, as shown in Sec. 3.3, HoP achieves a Lingo-Judge score of 67.8, establishing a new state-of-the-art. Ablation studies confirm that the three types of hints are complementary, collectively enhancing representation capability. Additionally, we explored various forms of Semantic and Affinity hints from different sources. Results indicate that as long as hint tokens represent token-wise affinity, performance improves, even when using a similarity matrix as Affinity hint tokens.
To summarize, we would like to highlight our contributions as follows:
  • We identify key limitations in current MLLMs combined with CLIP alone for VQA tasks in autonomous driving.
  • We introduce the Hints of Prompt (HoP) framework, which incorporates Affinity, Semantic, and Question hint tokens through a fusion module, significantly enhancing visual representation, enriching the LLM’s understanding of complex driving scenes, and boosting overall performance.
  • Extensive experiments demonstrate significant improvements of our proposed methods for VQA, establishing new state-of-the-art results on LingoQA [34], DRAMA [32], and BDD-X [21].
2 Related Work
Figure 3: A VQA sample case visualizes the impact of the three proposed hint types compared to the baseline.
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) [27, 26, 29, 8, 51, 49] have made significant advancements by integrating Large Language Models (LLMs) [10, 1, 43] with visual encoders, addressing diverse multimodal tasks across various domains.
In autonomous driving, MLLMs address critical challenges in perception, planning, control, question answering (QA), and generative world models. For example, GPT-Driver [33] reframes motion planning as a language modeling task, enhancing accuracy, while DriveGPT4 [48] combines vision and language inputs to predict control signals with real-time explanations, improving transparency.
In perception, MLLMs advance sensor data fusion from LiDAR and cameras; models like Talk2BEV [11] use Bird’s Eye View (BEV) features for complex visual tasks, while LMDrive [38] and DriveVLM [41] employ trajectory-based reasoning to refine scene understanding and decision-making. In QA, DriveLM [40] uses graph-based reasoning to enhance real-time planning. For world models, generative approaches like DriveDreamer [47] and GAIA-1 [19] generate realistic driving scenarios for data augmentation and training. These advancements highlight MLLMs’ potential to enhance multimodal understanding and optimize autonomous driving systems across diverse applications.
Enhancing Visual Representations in MLLMs. Effective alignment of visual representations with LLM capabilities is essential for improving overall performance. Some approaches focus on enhancing visual encoders directly. For instance, InternVL [8] introduces a large-scale visual encoder to improve feature granularity, while DeepSeek-VL [30] and Cambrian-1 [42] combine multiple encoders like SigLIP [50] and SAM [22] to leverage complementary strengths. To improve efficiency, BLIP-2 [23] employs a lightweight querying transformer to bridge vision and language models.
QA-ViT [17] incorporates textual information into visual features to help MLLMs focus on relevant image regions, enhancing interpretability while maintaining efficiency. Similarly, we utilize the same LLM text embeddings as Question hints, but our fusion strategy is more easily integrated with various visual backbones. Hint-AD [15] introduces a holistic alignment between natural language and intermediate perception and planning tokens, improving interpretability in complex scenarios. In contrast, we further introduce instance-level and question-level hint tokens with a lightweight fusion module, enabling the MLLM visual encoder to dynamically extract relevant hints for a better understanding of the current driving scenario.
3 Method
3.1 Overview
As shown in Fig. 1, the Hints of Prompt (HoP) framework leverages three distinct token types—Affinity, Semantic, and Question hints—each representing a different level of information. These hints progress from instance-level, token-wise connections in the Affinity hint, to high-level semantic information in the Semantic hint, and finally to question-specific context in the Question hint, gradually refining and aligning the visual representation with the LLM. A simple Hint Fusion module integrates these hints into a cohesive visual representation, which is then processed by an adapter and passed to the LLM for answer generation. By enabling the visual encoder to dynamically incorporate hints across different information levels, our approach enhances context-aware responses in VQA for autonomous driving scenarios.
3.2 Hint Tokens
Affinity hint The default practice of MLLMs [27, 30, 26] applies the CLIP [37] encoder for visual feature extraction, where the features are well-aligned with the text embedding space and contain rich semantic knowledge. However, the CLIP visual tokens struggle to understand driving scenes that require spatial reasoning among instances on the road. As shown in Fig. 2, we observe that the CLIP visual tokens drop the affinity information among instances, which might hinder the following LLM from reasoning upon these limited visual tokens. We find that DINOv2 [35] maintains significant affinity among tokens within the same instances, offering a more coherent spatial representation. Therefore, we use DINOv2 tokens as Affinity hint tokens to reinforce the instance-level structure in complex driving scenes.
To verify that the enhancement is due to token-wise affinity rather than specific DINOv2 features, we experiment with similarity information among the DINOv2 visual tokens and observe a notable performance gain. A detailed discussion is provided in Sec. 4.4.
Semantic hint After implementing instance-level enhancement, providing a semantic category specific to the driving domain can further focus visual representations on key instances, such as the cyclist shown in Fig. 3. The Semantic hint aims to incorporate high-level, domain-specific semantic information, such as vehicles, traffic signs, and pedestrians, which is crucial for understanding complex interactions. To achieve this, we employ sparse queries from DETR-like [6] detection or segmentation pipelines as Semantic hints. Specifically, GroundDINO [28] or Mask2Former [9] will be utilized for Semantic hint extraction. Experimental results (Tab. 5) show that, benefiting from Mask2Former’s adaptation to driving scenarios, it aligns more effectively with the task-specific requirements of VQA in autonomous driving.
Question hint In VQA tasks, different questions often shift attention to specific regions of an image. For example, a question about the weather may focus on broader aspects, such as the sky or overall lighting. In contrast, a question about pedestrians in the path directs attention to finer details and specific areas. Incorporating question-specific information can further refine this focus by enhancing visual representation to capture key instances in the driving environment. Aligning the visual representation with each question’s particular demands enables a more targeted, context-aware interpretation, thus enhancing relevance and effectiveness in complex driving scenarios. Similar to QA-ViT [17], we found that using LLM’s text tokens as the Question hint effectively achieves context-aware representation. We also explored an alternative approach that utilizes CLIP’s text encoder to generate the Question hint. A comparison of these methods is shown in Tab. 6.
As shown in Fig. 3, incorporating the three proposed hints significantly improves the model’s response over the baseline. The baseline incorrectly suggests overtaking is possible, overlooking oncoming traffic. With the added hints, the model gains instance-level structure and focuses more accurately on key elements, such as the cyclist and oncoming vehicles. This added context enhances the model’s spatial reasoning, resulting in a more precise and relevant response in complex driving scenarios. Detailed quantitative analysis is available in Sec. 4.
Figure 4: Different fusion strategies for the Hints Fusion module. The joint cross-attention strategy demonstrates the best performance. Residual connections are omitted for simplicity.
3.3 Hints Fusion
We explore several fusion strategies to merge the above hints into the original CLIP [37] visual tokens. Formally, we denote Affinity, Semantic, and Question hints as $\mathbf{H}_{\text{A}}\in\mathbb{R}^{N\times d}$, $\mathbf{H}_{\text{S}}\in\mathbb{R}^{M\times d}$, and $\mathbf{H}_{\text{Q}}\in\mathbb{R}^{K\times d}$, respectively, where $N$, $M$, and $K$ are the number of tokens for each hint, and $d$ is the embedding dimension. The original CLIP visual tokens are represented as $\mathbf{P}\in\mathbb{R}^{L\times d}$, and the fused visual tokens are denoted as $\mathbf{P}_{\text{f}}\in\mathbb{R}^{L\times d}$. The notation [·] indicates the concatenation operation, $\texttt{CA}(\mathbf{P},\mathbf{H})$ denotes a Cross-Attention layer with $\mathbf{P}$ as query and $\mathbf{H}$ as key-value, and SA(·) denotes a Self-Attention layer. As shown in Fig. 4, we present the different fusion strategies as follows.
Concatenation As a baseline strategy, hint tokens are directly concatenated with CLIP tokens, as illustrated in Fig. 4(a):
$\mathbf{P}_{\text{f}}=\texttt{[}\mathbf{P},\mathbf{H}_{\text{A}},\mathbf{H}_{\text{S}},\mathbf{H}_{\text{Q}}\texttt{]}.$ (1)
Interestingly, we find that simple concatenation can harm visual representation and degrade performance (see Tab. 7), suggesting that directly injecting multi-level information may confuse the adapter and LLM, hindering relevant information extraction.
Self-Cross-Attention To enable deeper interaction between visual tokens and hint tokens, we propose a Self-Cross-Attention module (Fig. 4(b)) that mimics the standard Transformer [44] layer but omits the feed-forward network (FFN) to accommodate the subsequent adapter module:
$\mathbf{P}_{\text{f}}=\mathbf{P}+\texttt{CA}(\texttt{SA}(\mathbf{P}),\texttt{[}\mathbf{H}_{\text{A}},\mathbf{H}_{\text{S}},\mathbf{H}_{\text{Q}}\texttt{]}).$ (2)
This allows refinement of the visual tokens before incorporating the hint tokens.
{booktabs}
colsep = 11pt, colspec = lccccc, row8-9 = bg=gray!15, cell12-6 = c, Method LLM Lingo-Judge $\uparrow$ BLEU-4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
GPT-4V [49] (zero-shot) - 59.6 6.30 12.4 42.8
LingoQA Baseline [34] Vicuna-v1.5-7B 60.6 15.0 18.6 65.6
QA-ViT [17]† Vicuna-v1.5-7B 63.6 14.6 19.0 65.9
VTS [31] InternLM2-7B 64.2 14.5 20.5 56.9
LLaVA-v1.5† Vicuna-v1.5-7B 63.2 14.1 19.3 63.7
LLaVA-v1.5 (DINOv2)† Vicuna-v1.5-7B 62.8 14.7 19.3 68.3
Efficient HoP Vicuna-v1.5-7B 66.8 15.2 20.0 66.2
HoP Vicuna-v1.5-7B 67.8 15.8 20.3 70.9
Table 1: Performance comparison on the LingoQA dataset. The best-performing method for each metric is highlighted in bold, while the second-best method is indicated by an underline. “†” indicates model that are fine-tuned on the LingoQA dataset. Efficient HoP represents the lightweight version of our HoP method, while LLaVA-v1.5 serves as our baseline. LLaVA-v1.5 (DINOv2) replaces the CLIP visual encoder with a DINOv2 encoder. The HoP method consistently demonstrates superior performance across all metrics.
Joint Cross-Attention As shown in Fig. 4(c), inspired by Flamingo [2], Joint Cross-Attention omits the self-attention stage and uses the visual representation $\mathbf{P}$ as both query and key-value in a single cross-attention layer:
$\mathbf{P}_{\text{f}}=\mathbf{P}+\texttt{CA}(\mathbf{P},\texttt{[}\mathbf{P},\mathbf{H}_{\text{A}},\mathbf{H}_{\text{S}},\mathbf{H}_{\text{Q}}\texttt{]}).$ (3)
The design simplifies the fusion process without weakening the interaction between visual tokens and hint tokens.
To compare the effects of different query-hint interaction strategies, we further explore two additional approaches: Parallel Cross-Attention and Sequential Cross-Attention, as shown in Fig. 4(d, e). In these strategies, hint tokens first interact individually with the query and are then fused through either a parallel or cascading way.
Sequential Cross-Attention sequentially applies cross-attention between the visual representation and each hint, progressively enriching the representation:
$\mathbf{P}_{\text{f}}=\mathbf{P}+\texttt{CA}(\texttt{CA}(\texttt{CA}(\mathbf{P},\mathbf{H}_{\text{A}}),\mathbf{H}_{\text{S}}),\mathbf{H}_{\text{Q}}).$ (4)
Parallel Cross-Attention applies cross-attention simultaneously between the visual representation and each hint token, which allows all hint tokens to influence the visual representation in parallel:
$\mathbf{P}_{\text{f}}=\mathbf{P}+\texttt{CA}(\mathbf{P},\mathbf{H}_{\text{A}})+\texttt{CA}(\mathbf{P},\mathbf{H}_{\text{S}})+\texttt{CA}(\mathbf{P},\mathbf{H}_{\text{Q}}).$ (5)
We empirically find that the proposed Joint Cross-Attention achieves the best performance with high efficiency across diverse VQA tasks in autonomous driving, outperforming the other fusion strategies mentioned above. More details are provided in Sec. 4.4.
3.4 Efficient Hints of Prompt
Although introducing extra hint tokens significantly improves VQA performance, extracting these hints adds a computational burden. To reduce resource overhead, we propose an efficient version of HoP. The main idea is to distill or fine-tune lightweight heads to mimic the behavior of DINOv2 [35] or Mask2Former [9], providing additional instance-level and semantic-level hints to enhance the visual representation while maintaining low latency. More details are provided in Sec. 4.1 and Sec. 4.4.
4 Experiments
We evaluate the Hints of Prompt (HoP) across three datasets: LingoQA [34], DRAMA [32], and BDD-X [21], focusing on its performance in VQA tasks for autonomous driving. Our experiments encompass overall performance improvements, along with ablation studies on the design of each hint, fusion strategies, and efficiency.
{booktabs}
[…]
LCP [32] 73.9 54.7 39.1 70.0 3.7 56.0
VTS [31] 75.3 55.8 40.7 74.7 2.8 58.0
LLaVA-v1.5† 75.8 56.1 41.0 78.0 2.9 58.4
Efficient HoP 76.0 56.2 41.3 78.5 2.7 58.8
HoP 76.2 56.3 41.7 79.8 2.8 59.1
Table 2: Performance comparison on the DRAMA dataset. “†” indicates model that are fine-tuned on the DRAMA dataset. HoP demonstrates enhanced performance over other methods across most metrics, with LLaVA-v1.5 as the baseline.
{booktabs}
[…]
ADAPT [20] 100.93 20.90 46.17 62.66 16.44 40.80 52.71 13.56 40.49 85.38 17.40 43.04
DriveGPT4 [48] 113.20 20.38 46.46 65.01 16.94 40.51 57.29 12.28 42.07 99.10 18.32 44.73
LLaVA-v1.5† 111.84 19.83 47.57 63.76 14.69 40.34 49.80 11.60 41.97 96.51 17.42 45.42
Efficient HoP 118.30 19.95 47.65 64.80 17.58 40.73 58.20 12.92 42.58 99.80 17.70 45.65
HoP 121.01 20.09 47.60 65.16 18.02 40.81 60.37 13.01 43.52 102.20 17.91 45.79
Table 3: Performance comparison across different splits of the BDD-X testing dataset. “†” indicates model that are fine-tuned on the BDD-X dataset. The evaluation follows the DriveGPT4 [48] protocol, reporting CIDEr, BLEU-4 (B4), and ROUGE scores across Easy, Medium, Hard, and Overall (All) splits. LLaVA-v1.5 serves as the baseline.
4.1 Implementation Details
We implement the Hints of Prompt (HoP) method based on LLaVA-v1.5 [26], freezing the visual encoder and fully fine-tuning the remaining modules.
For the Affinity hint, we use DINOv2-large [35] to extract 576 hint tokens by downsampling the feature map to align with the CLIP tokens ($24\times 24$). To verify that the improvement is due to token-wise affinity rather than specific DINOv2 features, we compute a similarity matrix from the 576 DINOv2 tokens and use it as hint tokens, processed through a linear layer. Results are shown in Tab. 4.
For the Semantic hint, we select the top-K queries from Mask2Former [9] or GroundingDINO [28] as hint tokens, based on confidence scores, with each token adding an embedding feature that corresponds to the label index. The Question hint uses the LLM’s textual embeddings associated with question tokens.
To create Efficient HoP, we distill and fine-tune two lightweight heads from the CLIP backbone: a 4-layer, 8-heads, 512-dimensional ViT-like [16] decoder for the Affinity hint, distilled via cosine similarity to DINOv2 features on the LingoQA training dataset; and a ViTDet-like [24] neck with a Mask2Former head for the Semantic hint, trained on the Cityscapes dataset [13]. The training setup follows that of the original Mask2Former.
The whole HoP training is conducted for one epoch using the AdamW optimizer with a cosine learning rate schedule, a warm-up ratio of 0.03, and a base learning rate of 2e-5. All models are trained on 32 NVIDIA A100 GPUs with a batch size of 4 per GPU.
Figure 5: Impact of each hint type within the HoP framework. Abbreviations: AH (Affinity hint), SH (Semantic hint), QH (Question hint). “w/ cls” indicates the addition of class information to the Semantic hint, and “ALL” denotes the combination of all three hint types.
4.2 Datasets and Benchmarks
LingoQA [34] is a benchmark dataset for video question answering in autonomous driving, with 28K unique driving scenarios and 419K question-answer (QA) pairs across tasks like object recognition, scene description, driving reasoning, and action justification. The dataset consists of two subsets: Action and Scenery, plus a high-quality Evaluation subset for benchmarking. The main metric, Lingo-Judge, is an evaluation model aligned with human preferences, while BLEU [36], METEOR [5], and CIDEr [45] provide lexical similarity assessments.
DRAMA [32] dataset comprises 17K video clips of interactive urban traffic scenes in Tokyo, and provides extensive annotations for autonomous driving risk assessment. It includes lots of VQA pairs about 11.8K training samples and 2.5K testing samples. Evaluation metrics include BLEU, METEOR, CIDEr, ROUGE [25], and SPICE [3].
BDD-X [21] is a dataset developed for autonomous driving tasks, focusing on video captioning and rationales, but it lacks native VQA pairs. To address this, we adopt the VQA pairs and evaluation protocol provided by DriveGPT4 [48], which include 26K training samples and 1.8K testing samples. Metrics such as BLEU, CIDEr, and ROUGE are used to ensure consistency and comparability of results.
4.3 Comparison with Existing Works
Our results on the LingoQA [34], DRAMA [32], and BDD-X [21] datasets demonstrate the effectiveness of the Hints of Prompt (HoP) method in enhancing VQA for autonomous driving. As shown in Sec. 3.3, HoP significantly outperforms baseline models on LingoQA, achieving a top Lingo-Judge score of 67.8 and leading BLEU-4, METEOR, and CIDEr scores, reflecting strong semantic accuracy and alignment with ground truth answers.
As shown in Sec. 4, HoP surpasses other methods on the DRAMA dataset across key metrics, including BLEU-1, BLEU-4, METEOR, ROUGE, and SPICE, highlighting its capability to generate detailed and contextually accurate answers in urban traffic scenes.
In Sec. 4, HoP’s performance on BDD-X, evaluated using the DriveGPT4 protocol, consistently outperforms other methods across all difficulty levels (Easy, Medium, Hard, and All), achieving the highest CIDEr scores and robust results even in challenging driving scenarios.
In addition, the Efficient HoP, a lightweight version of our method, also performs competitively in all datasets, demonstrating the scalability of our approach.
Visualization examples can be found in Fig. 7.
AH types LJ $\uparrow$ B4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
Original tokens 64.6 14.6 19.0 66.0
Similarity tokens 64.2 15.5 19.3 71.1
Table 4: Impact of different Affinity hint types on performance. This ablation examines only the Affinity hint without incorporating other hint types. “Original tokens” refer to DINOv2 tokens, while “Similarity tokens” are derived from the similarity matrix. Abbreviations: AH (Affinity Hint), LJ (Lingo-Judge), B4 (BLEU-4).
SH types LJ $\uparrow$ B4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
GroundingDINO [28] 66.4 15.2 20.1 68.4
Mask2Former [9] 67.8 15.8 20.3 70.9
Table 5: Impact of different Semantic hint types on performance. Abbreviations: SH (Semantic hint), LJ (Lingo-Judge), B4 (BLEU-4).
QH types LJ $\uparrow$ B4 $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
CLIP 65.2 15.3 20.1 67.0
LLM 67.8 15.8 20.3 70.9
Table 6: Impact of different Question hint types on performance. “CLIP” refers to hint tokens derived from the CLIP text encoder, while “LLM” refers to tokens from the LLM input embedding layer. Abbreviations: QH (Question Hint), LJ (Lingo-Judge), B4 (BLEU-4).
4.4 Ablation Study
To evaluate the contributions of each component within the HoP framework, we conduct comprehensive ablation studies. These include analyzing the effectiveness of each hint type independently (Fig. 5), assessing the impact of different hint sources (Tab. 4, Tab. 5, Tab. 6), and hyperparameters (Fig. 6). We also evaluate various fusion strategies for integrating the hints (Tab. 7) and, finally, analyze the efficiency gains of the proposed Efficient HoP model (Tab. 8).
Combination of Different Hint Types Fig. 5 shows an ablation study on the LingoQA dataset, assessing the impact of each hint type within the HoP framework. Applying each hint independently (IDs 1–3) consistently improves performance over the baseline (ID 0) across all metrics, highlighting each hint’s unique contribution to model comprehension. Combining two hints (IDs 4–6) provides additional gains, while using all three hints together (IDs 7–8) achieves the highest scores overall, demonstrating their complementary effects. Notably, adding class information to the Semantic hint (+cls in ID 8) further enhances performance.
Figure 6: The effect of the number of Semantic hint tokens on performance in the LingoQA [34] dataset.
Effectiveness of Affinity Hint As shown in Table 4, using DINOv2 [35] token-wise similarity as the source for the Affinity hint provides notable improvements over the baseline, indicating that the core benefit of the Affinity hint lies in capturing token-wise similarity. Although using the full DINOv2 features as the Affinity hint achieves slightly better results, this can be attributed to the additional feature information present in DINOv2. These results confirm the effectiveness of the Affinity hint and underscore the importance of token-wise connections in representing driving scenarios.
Impact of Different Semantic Hint As shown in Tab. 5, Mask2Former [9] achieves superior performance across all metrics, with a Lingo-Judge score of 67.8 compared to 66.4 for GroundingDINO [28]. This boost suggests that Mask2Former’s domain adaptation through training on Cityscapes [13] allows it to capture driving-specific semantic details more effectively. In contrast, while GroundingDINO excels in open-set object detection, its lack of domain-specific training limits its alignment with the task-specific requirements of VQA in autonomous driving.
We also analyze the impact of varying the number of Semantic hint tokens on performance in the LingoQA dataset. As shown in Fig. 6, increasing the token count from 8 to 16 improves performance across all metrics. Adding more tokens up to 32 yields similar results, while increasing the count to 64 leads to a decline, likely due to noise introduced by less relevant queries from Mask2Former.
Figure 7: Visualization results across datasets. Example responses generated by our method on the LingoQA [34], DRAMA [32], and BDD-X [21] datasets.
Impact of Different Question Hint Sources We evaluate two sources for extracting question embeddings for the Question hint: CLIP’s text model and the LLM embedding layer. As shown in Tab. 6, using the LLM embedding layer yields higher performance across all metrics, achieving a Lingo-Judge score of 67.8 compared to 65.2 with CLIP’s text model. This result suggests that the LLM embeddings offer superior alignment for multimodal reasoning tasks.
Fusion Strategy LJ $\uparrow$ #Params (M) GFLOPs
Concatenation 58.8 - -
Self-Cross-Attention 65.2 12.9 5.0
Joint Cross-Attention 67.8 8.7 3.8
Sequential Cross-Attention 67.0 17.1 8.63
Parallel Cross-Attention 66.6 17.1 8.63
Table 7: Impact of different fusion strategies for integrating hints. Joint Cross-Attention delivers the best performance with efficient parameter and computational use.
Method #Params (M) Latency (ms)
LLaVA-v1.5 303.5 1535
HoP 832.0 2251
Efficient HoP 385.8 1548
Table 8: Efficiency of the Efficient HoP. #Params indicates the parameter count for all modules related to visual representation. Latency reports the average inference time measured on the LingoQA [34] validation set on an NVIDIA RTX 4090 GPU.
Hint Fusion Strategy As shown in Tab. 7, the Joint Cross-Attention strategy delivers the best performance, achieving the highest Lingo-Judge (LJ) score of 67.8. This method is not only effective but also highly efficient, requiring only 8.7M parameters and 3.8 GFLOPs—significantly less than the Self-Cross-Attention. In comparison, Sequential and Parallel Cross-Attention strategies, while competitive with LJ scores of 67.0 and 66.6, respectively, require nearly twice the computational resources of Joint Cross-Attention.
Efficient HoP In the original HoP framework, we use DINOv2 [35] and Mask2Former [9] for Affinity and Semantic hints, which boost performance but increase computational costs. To reduce these, we create an Efficient HoP. As shown in Tab. 8, Efficient HoP introduces fewer additional parameters and latency compared to the HoP. Moreover, as shown in Sec. 3.3, Sec. 4, and Sec. 4, Efficient HoP surpasses the baseline across various metrics and achieves results close to those of the full HoP model.
5 Conclusion
In this paper, we introduce the Hints of Prompt (HoP) framework to address key limitations of current MLLMs in VQA tasks for autonomous driving. By incorporating Affinity, Semantic, and Question hints, HoP effectively captures instance-level structure, domain-specific elements, and query-specific context, leading to enhanced interpretability and reasoning capabilities. Extensive experiments on the LingoQA, DRAMA, and BDD-X datasets demonstrate that HoP consistently achieves state-of-the-art performance, validating the effectiveness of our proposed method. Additionally, we propose an Efficient HoP variant that reduces computational overhead while maintaining high performance. We believe that HoP’s structured integration of multi-level hints offers a promising direction for advancing VQA in complex, safety-critical scenarios.